Allison drove home at 70 mph, but her brother Austin, who left at the same time, could drive at only 54 mph. When Allison arrived, Austin still had 48 miles to go. How far did Allison drive?
Found 2 solutions by ikleyn, solver91311:
Answer by ikleyn(52824)

.
Let t be the time they drove.


Then you have this "distance" equation


    70*t = 54*t + 48,


saying that both parts of the equation represent the same distance.  Then


    70t - 54t = 48,


    16t = 48  ====>  t = 48/16 = 3 hours.


Hence, the distance is  3*70 = 210 miles.

Solved.

There are other ways to solve the problem, but this one is the simplest, probably.

---------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.


Answer by solver91311(24713)



Distance equals rate times time, so Allison's trip is modeled by while Austin's trip is modeled by . We can use the same in each model because the time of travel is given as the same.

But we know that , so



Which is to say:



But that means



Solve for , then calculate


John

My calculator said it, I believe it, that settles it